OpenStudy (anonymous):

find solutions for 9x^2-81=0

OpenStudy (anonymous):

x = 3 or -3

OpenStudy (anonymous):

x^2=9 and so u get x=+3 and x=-3

OpenStudy (anonymous):

x^2-9 (x+3)(x-3)=0

OpenStudy (anonymous):

(3x+9)(3x-9) so x = -9 or 9

OpenStudy (anonymous):

what about for x^2-9=0, will that also be -3 and3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use the difference of 2 squares why? 9x^2 is a square and 81 is a square difference = - sqr root 9x^2 =3x sqr root 81=9 answers are solutions to (3x+9)(3x-9)=0 3x+9=0 implies 3x=-9 implies x=-3 and 3x-9 implies 3x=9 implies x=3 answers as given earlier x=3 or -3

OpenStudy (anonymous):

x^2-9=0 sqr root x^2 is x sqr root of 9 =3 0=(x+3)(x-3) x=3 or -3 yes the answers are the same.